OpenStudy (anonymous):

Classify as either a quadratic equation, a linear equation, or neither. f(x) = x^2 + 2x + 3 Quadratic Neither Linear

OpenStudy (princeharryyy):

every quadratic has one of the power of x as x^2

OpenStudy (princeharryyy):

quad ans A

OpenStudy (anonymous):

@eviI

OpenStudy (princeharryyy):

no powers above 2 are valid in case of quadratic equations @keridavis

OpenStudy (anonymous):

@RedNeckOutLaw

OpenStudy (princeharryyy):

any function of the form y= ax^2 + bx +c is known as quadratic equation

OpenStudy (anonymous):

So it's a??

OpenStudy (princeharryyy):

why are u getting confused @keridavis

OpenStudy (anonymous):

I have no idea. I'm lost

OpenStudy (princeharryyy):

yeah !!!!

OpenStudy (princeharryyy):

where are u getting lost

OpenStudy (anonymous):

linear equations are usually in this form: y = mx + b quadratics: y = ax^2 + bx + c exponential: y = a^x

OpenStudy (anonymous):

Wouldn't it be linear???

OpenStudy (anonymous):

of exponetial/

OpenStudy (princeharryyy):

linear function is an equation of a line not any curve y= mx + c is an equation of line and linear originate from line , x has power 1

OpenStudy (anonymous):

@zepdrix

OpenStudy (princeharryyy):

for any function to be an exponential a constant number has its power as x example, y = 2^x , y= 5 ^x , y = a^x where a is some constant like this @keridavis

OpenStudy (anonymous):

i still think linear...

OpenStudy (princeharryyy):

i told you what linear means linear means a line and equation of line is y = ax +b

OpenStudy (anonymous):

yes so it shoudl be because it's x=x+4

OpenStudy (princeharryyy):

while equations like y = ax^2 + bx +c are quadratic equations

OpenStudy (princeharryyy):

what is x = x+4

OpenStudy (princeharryyy):

f(x) equals the value of y

OpenStudy (princeharryyy):

@keridavis

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (princeharryyy):

and f(x) = x^2 + 2x + 3 => a=1, b=2, c=3 , satisfy it in f(x) = ax^2 + bx + c which means quadratic equation

OpenStudy (princeharryyy):

@keridavis

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(x) = x^2 + 2x + 3}\) is just like ax²+bx+c So which option do you pick ?

OpenStudy (anonymous):

I say Linear

OpenStudy (princeharryyy):

and f(x) = x^2 + 2x + 3 => a=1, b=2, c=3 , satisfy it in f(x) = ax^2 + bx + c which means quadratic equation @keridavis

OpenStudy (princeharryyy):

do you know little bit about graphs and curves @keridavis

OpenStudy (anonymous):

@dan815

OpenStudy (solomonzelman):

I am going to show you something

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(x) = x^2 + 2x + 3}\) \(\normalsize\color{blue}{ f(x) = x^2 + 2x +1 + 2}\) \(\normalsize\color{blue}{ f(x) =(x+1)^2 + 2}\)

OpenStudy (solomonzelman):

Parabola centered at \(\normalsize\color{blue}{ (-1,2)}\) opening up.

OpenStudy (princeharryyy):

take the feel of quadratic @ keridavis what do you think this is ?

OpenStudy (anonymous):

This isn't helping!!!!!

OpenStudy (princeharryyy):

take a look at this figure @keridavis what can u figure out of this image///

OpenStudy (anonymous):

@SolomonZelman please help me on this..

OpenStudy (princeharryyy):

@keridavis why aren't u getting .... do u know what linear means

OpenStudy (anonymous):

Is ist linear or quad???

OpenStudy (princeharryyy):

quadratic 101 %